OpenStudy (anonymous):

LAST QUESTION! A long one.. Given: 7x + 2 = 2 + 5x – 6 Prove: x = –3

OpenStudy (anonymous):

Given the equation 7x + 2 = 2 + 5x – 6, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 7x + 2 = 2 – 6 + 5x. Then, use the ____________________ to group the like terms. This gives the equation 7x + 2 = (2 – 6) + 5x. Next, combine like terms to get the equation 7x + 2 = –4 + 5x. Use the subtraction property of equality to subtract 5x from both sides of the equation. This gives the equation 2x + 2 = –4. Then use the subtraction property of equality again to subtract 2 from both sides of the equation. This gives the equation 2x = –6. Finally, use the division property of equality to divide both sides of the equation by 2 to give a final solution of x = –3. Therefore, given the equation 7x + 2 = 2 + 5x – 6, x is equal to –3.

OpenStudy (anonymous):

Which justification was left out of the paragraph proof above? Answer Commutative Property of Addition Associative Property of Addition Distributive Property Addition Property of Equality

OpenStudy (lgbasallote):

"group the terms" that's kinda a giveaway isnt it?

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

Commutative property: When two numbers are added, the sum is the same regardless of the order of the addends. For example 4 + 2 = 2 + 4 Associative Property: When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example (2 + 3) + 4 = 2 + (3 + 4) Additive Identity Property: The sum of any number and zero is the original number. For example 5 + 0 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3 CREDIT: http://www.aaastudy.com/add74ax1.htm

OpenStudy (anonymous):

so its associative property

OpenStudy (lgbasallote):

Commutative property -> a + b = b + a Associative property (a+b) + c = a + (b+c) Distributive property -> a(b+ c) = ab + ac Additive Property -> a - b = c -> a = c + b

OpenStudy (anonymous):

lol i was about to say B xD. Ty. Igba and iHelp. I think i understand the rules a little better now.

OpenStudy (lgbasallote):

well it is B :p

OpenStudy (anonymous):

:) ty for your help.